OpenStudy (anonymous):

The hypotenuse of a right triangle is 25 cm, and the shorter leg is 15 cm. Find the length of the other leg. 2√5cm?

OpenStudy (amistre64):

leg^2 + leg^2 = slant^2 leg^2 = slant^2 - leg^2 leg = sqrt(slant^2 - leg^2 )

OpenStudy (amistre64):

leg = sqrt(25^2-15^2) leg = sqrt(625-225) leg = sqrt(400) leg = sqrt(4*100)

OpenStudy (anonymous):

20cm?

OpenStudy (amistre64):

thats better